find the center and radius given the following equation of a circle
x^2+y^2+6x+2y+6=0

Answer by dabanfield(803)

You need to get the equation into the form:
(x-h)^2 = (y-k)^2 = r^2
This is the equation of a circle with center (h,k) and radius r.
Reorganizing the terms given we have:
(x^2 + 6x) + (y^2 + 2y) = - 6
"Complete the squares" for the x and y terms:
(x^2 + 6x + 9)+ (y^2 + 2y + 1) = -6 +9 + 1 = 4
Factoring the terms:
(x+3)^2 + (y+1)^2 = 4
Now the equation is in the form needed to find the center and the radius.